A firm has the production function f(K,L)=4KL^1/2. The firm is currently using 16 units of labor and 2 units of capital. What is the firm's marginal product of capital?
Answer #1

Marginal product of capital (MPk) = dQ / dK = d[f(K,L)] / dK = (4L^1/2* 1)+(4K*0) = 4L^1/2

If L = 16, MPk = 4(16)1/2 = 4*4 = 16